# Color Equivariant Network

### v2

@@ 5 Experiments
$A$/$A$ $A$/$B$ Params
Z2CNN 1.54 (0.10) 57.38 (22.06) 22,130
Hue-4* 1.97 (0.25) 2.08 (0.15) 25,690
Hue-3* 1.79 (0.25) 1.81 (0.28) 22,658
CEConv 1.79 (0.13) 1.83 (0.19) 28,739
CEConv-2 2.81 (0.44) 5.08 (0.18) 30,539
Table 1: Classification error on the hue shift MNIST dataset. Our networks are marked with (*). The number of parameters for all networks is reported.
In this section we demonstrate the strength of our method in the presence of global and local hue variation, and global saturation variation on synthetic and real world datasets. We also demonstrate the ability of our model to perform color dependent tasks such as sorting.
(a) Z2CNN      (b) Hue-CNN      (c) Hybrid
Figure 4: Comparison of feature maps for equivariant and non-equivariant architectures. In the conventional network (a), the relationship between feature maps is unclear. In our model (b), a hue shift in the input results in a predictable hue shift of the internal representations. In the hybrid network (c), the feature maps appear approximately invariant to hue shift.
5.1 Hue-shift MNIST: Global hue equivariance
We evaluate the performance of our models in the presence of global hue shifts on a classification dataset derived from MNIST [25]. The dataset consists of 70k examples, 60k of which are used for training and 10k of which are used for testing.
To test the performance of our models, we partition the testing dataset into two subsets: an in-distribution testing set and an out-of-distribution testing set. In the training set and in-distribution testing set, denoted subset $A$, digits are colored with a randomly selected hue between $0^{\circ}$ and $240^{\circ}$ (see Figure 6(a)). In the out-of-distribution testing set, subset $B$, digits are colored with a randomly selected hue between $240^{\circ}$ and $360^{\circ}$ (see Figure 6(b)).
Network Group $A$/$A$ $A$/$B$ $A$/$C$ Params
ResNet18 $\mathbb{Z}^{2}$ 0.00 (0.00) 38.53 (3.19) 17.73 (0.90) 11.17M
Hue-4* 0.00 0.00 0.00 11.17M
ResNet44 $\mathbb{Z}^{2}$ 0.00 (0.00) 41.34 (1.73) 22.29 (10.80) 2.64M
Hue-4* 0.00 (0.00) 0.00 (0.00) 0.00 (0.00) 2.63M
CNN $\mathbb{Z}^{2}$ 0.00 (0.00) 51.25 (9.59) 26.66 (19.60) 20,192
@@ 5.1 Hue-shift MNIST: Global hue equivariance
Hue-3* 0.00 (0.00) 0.03 (0.04) 0.04 (0.03) 19,478
CEConv 0.00 (0.00) 0.02 (0.03) 0.05 (0.04) 26,441
Table 2: Classification error on the 3D Shapes dataset. Our networks are marked with (*). The number of parameters for all networks is reported.
Model performance is reported in Table 1. On the in-distribution test case ($A$/$A$), the performance of our hue-equivariant models (Hue-x) and the conventional model (Z2CNN) are comparable. On the out-of-distribution test case ($A$/$B$), the performance of our hue-equivariant models is preserved, while the performance of the conventional model deteriorates significantly.
The performance gap on the out-of-distribution test case can be attributed to the difference in internal representation structure. To understand this better, we generate feature representation trajectories in the penultimate layer of our network and the baseline architecture. Feature representation trajectories are generated by continuously varying the hue of an input image, and visualized using tSNE projection (see Figure 3). For the baseline architecture, the trajectories of different digits overlap for some hues (Figure 3(a)). For our model, the trajectories of different digits are confined to separate clusters (Figure 3(b)).
We also compare our model to the hue-equivariant (CEConv), and hybrid network (CEConv-2) proposed in [29]. With the same hue discretization (i.e., $N=3$), our model performs comparably with the baseline hue-equivariant model, and out performs the baseline hybrid model. We investigate differences in the internal representations of our model, the conventional architecture, and the baseline hybrid model in Figure 4. For each network, we visualize the feature maps generated for a hue rotated input. For a hue discretization of $N=3$, the network feature maps have three channels and can be visualized as an RGB image. Only the feature maps of our network shift predictably in response to hue shifts of the input.
5.2 Hue-shift 3D Shapes: Local hue equivariance
We evaluate the performance of our models in the presence of local hue shifts on the 3D Shapes classification dataset [5]. The dataset consists of RGB images of 3D shapes; the color of the shape, the floor, and the walls vary across examples, as do the scale and orientation of the shape.
To test the performance of our models, we partition the dataset into four subsets: a training set of 48k examples, and a test set of 12k examples. In-distribution and out-of-distribution test sets differ by color range assigned to the shape, floor and walls. In the training set and in-distribution testing set (subset $A$), the color of the wall, the floor and the shape are randomly selected from the first half of the color space (colors 0-4 as defined in [5]). The out-of-distribution testing sets are designed to measure robustness to global hue shift, and local hue shift. In the out-of-distribution testing set designed to measure robustness to global hue shift (subset $B$), the color of the wall, the floor and the shape are randomly selected from the second half of the color space (colors 5-9 as defined in [5]). In the out-of-distribution testing set designed to measure robustness to local hue shift (subset $C$), the color of the wall and the floor are randomly selected from the first half of the color space, and the color of the shape is randomly selected from the second half of the color space.
Model performance is reported in Table 2. Our hue-equivariant architectures outperform all baseline models for all test cases. Noteably, our hue-equivariant architecture with CNN backbone and $\sim$22k parameters outperforms both the ResNet44 and ResNet18 architectures with $\sim$2M and $\sim$11M parameters respectively.
Error Params
@@ 5.2 Hue-shift 3D Shapes: Local hue equivariance
Sat-$d_{1}$* 19.60 (2.08) 23.3M
Sat-$d_{2}$* 16.08 (2.68) 23.3M
Sat-$d_{3}$* 24.53 (7.13) 23.3M
Sat-$d_{4}$* 23.57 (6.04) 23.3M
Table 3: Classification error on the Camelyon17 dataset. Our networks are marked with (*). The number of parameters for all networks is reported.
Figure 5: CIFAR-10 automobile class sorted by hue.
5.3 Camelyon17: Saturation shift in the wild
We evaluate the performance of our model in the presence of natural saturation shifts on the Camelyon17 classification dataset [1]. The dataset consist of images of human tissue collected from five different hospitals. Variation in tissue images results from variation in the data collection and processing procedures. The dataset consists of 387,490 examples, 302,436 of which are used for training and 85,054 of which are used for testing. Hospitals 1, 2, and 3 are represented in the training set, hospital 5 is represented in the testing set.
We present experimental results for four discretizations of the saturation space. Our choice of discretizations, i.e., $d\in\{\nicefrac{{1}}{{20}},\nicefrac{{1}}{{10}},\nicefrac{{3}}{{20}},\nicefrac{{1}}{{2}}\}$, are determined with consideration of the saturation range of the training set. For a given choice of discretization $d$, an element $s_{i}\in S_{N}$ shifts the saturation of an input image by $s_{i}=i*d*255$. We limit computational expense of lifting and convolution by performing these operations over the truncated set $\{s_{-1},s_{0},s_{1}\}$. Model performance is reported in Table 3. The performance of our saturation-equivariant models out perform all other models, and the performance of our hue-equivariant model is on par with the baseline model.
5.4 CIFAR-10: Color shift in the wild
We evaluate the performance of our model in the presence of natural hue shifts on the CIFAR-10 [23] classification dataset. The dataset consists of 60k examples, 40k of which are used for training, 10k for validation, and 10k for testing. Model performance is reported in Table 4. The performance of our color-equivariant models are comparable to the baseline model.
@@ 5.4 CIFAR-10: Color shift in the wild
While our hue-equivariant model performs on par with the baseline model, the structure of the representation can be leveraged for tasks that are not possible with the baseline alone. To illustrate this, we use our model to sort images in the automobile class by hue (see Figure 5). To produce an hue ordering, we compute the hue shifted pairwise Euclidean distance between the penultimate layer hue group feature maps of instances in the automobile class. Images $x_{1}$ and $x_{2}$ are close in the hue space if
$d(\Phi(x_{1}),\Phi(x_{2}))=\min_{i}d(h_{i}\Phi(x_{1}),\Phi(x_{2})),$ (20)
@@
where $\Phi(x_{i})$ denotes the penultimate layer hue group feature map for image $i$.
Error Params
ResNet44 7.86 (1.14) 2.64M
Hue-4* 8.83 (0.64) 2.63M
Sat-$d_{1}$* 10.45 (1.76) 2.55M
Sat-$d_{2}$* 10.26 (0.43) 2.55M
Sat-$d_{3}$* 9.24 (0.27) 2.55M
Hue-Sat-$d_{1}$* 10.68 (0.78) 2.50M
Hue-Sat-$d_{2}$* 11.13 (0.84) 2.50M
Hue-Sat-$d_{3}$* 11.16 (0.38) 2.50M
Hue-3* 8.42 (0.39) 2.55M
CEConv 8.86 (0.33) 1.15M
Table 4: Classification error on the CIFAR-10 dataset. Our networks are marked with (*). The number of parameters for all networks is reported.
@@ 6 Conclusion
In this paper we address the challenge of learning hue- and saturation-equivariant representations. Leveraging the observation that these perceptual transformations have geometric structure, we propose a group structure for each, and a group convolutional neural network that is equivariant to the transformations by design. By encoding perceptual variation in a finite group structure, we can use learned representations for color based sorting. Moreover, with an efficient implementation, our network achieves comparable performance to conventional architectures with comparable parameters.
@@ 6.1 Limitations and future work
While our method is equivariant to hue and saturation by design, it is not guaranteed to be equivariant to luminance. However, since luminance is computed as a linear combination of the red, green and blue channel values of an input image, if the preprocessing statistics (e.g., data mean) on the training set are valid for the testing set, luminance is approximately unchanged.
Additional limitations derive from our choice of equivariant modeling structure. Group equivariant networks are more computationally expensive than their conventional counterparts since they require computation of the filter orbit at each layer. This is compounded with the requirement for a finite filter orbit which can at best approximate a continuous group. Future work will investigate a Fourier series representation of the group to resolve both of the aforementioned challenges.
7 Acknowledgements
We would like to thank members of the CAB Lab for valuable discussions and feedback, and the Princeton Department of Computer Science for providing compute resources. This work is partially supported by the Princeton School of Engineering and Applied Science.
@@ 7 Acknowledgements
A Datasets
A.1 Hue-shift MNIST
We introduce the Hue-shift MNIST dataset to evaluate the performance of our equivariant architectures in the presence of global hue shift. We show select examples from the in-distribution and out-of distribution datasets in Figures 6(a) and 6(b).
(a) Dataset $A$ examples
(b) Dataset $B$ examples
A.2 Hue-shift 3D Shapes
We introduce the Hue-shift 3D Shapes dataset to evaluate the performance of our equivariant architectures in the presence of local hue shift. We show select examples from the in-distribution and out-of distribution datasets in Figures 7(a)- 7(c).
(a) Dataset $A$ examples
(b) Dataset $B$ examples
(c) Dataset $C$ examples
B Network architecture and training details
B.1 Hue-shift MNIST
We train our equivariant networks and the conventional architectures for 5 epochs with a batch size of 128. We optimize over a cross-entropy loss using the Adam optimizer [21] with $\beta_{1}=0.9$, and $\beta_{2}=0.999$. We use an initial learning rate of $10^{-3}$ for the $\mathbb{Z}^{2}$ network, and $10^{-4}$ for the hue-equivariant network. We train CEConv architectures using the hyperparameters reported in [29].
B.2 Hue-shift 3D Shapes: Local hue equivariance
For this task we compare against the ResNet18 architecture proposed in [19], and the Z2CNN and ResNet44 architectures proposed in [7]. Our hue-equivariant architectures are designed with the same network structure as the baselines, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our hue-equivariant architectures, we perform hue group pooling to yield an hue-invariant representation.
Following [34], we train the ResNet architectures for 100k iterations with a batch size of 128. We optimize over a cross-entropy loss using SGD with a learning rate of $10^{-2}$. Images are down-sampled by a factor of 2 to train the CNN architectures. CNN architectures are trained as described in Section B.1.
B.3 CIFAR-10: Color shift in the wild
For this task, we compare against the ResNet44 architecture proposed in [7]. Our hue-, saturation- and hue-saturation-equivariant architectures have the same number of layers as ResNet44, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 128. We optimize over a cross-entropy loss using SGD with an initial learning rate of $10^{-1}$ and a cosine-annealing scheduler. We train CEConv architectures using the hyperparameters reported in [29].
B.4 Camelyon17: Saturation shift in the wild
For this task, we compare against the ResNet50 architecture [19]. Our hue- and saturation-equivariant architectures have the same number of layers as ResNet50, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our hue- and saturation-equivariant architectures, we perform a group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 10k iterations with a batch size of 32. We optimize over a cross-entropy loss using the Adam optimizer [21] with an initial learning rate of $10^{-2}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. We train CEConv architectures using the hyperparameters reported in [29].
C Hue and saturation group action
@@ C.1 Hue group action
C.2 Saturation group action

### v3

@@ 5 Experiments
In this section, we highlight improved sample efficiency over a conventional CNN baseline, improved hue-equivariance over CEConv [16], the utility of saturation equivariance on the Camelyon17 dataset [31], and the utility of our architectures for color based sorting. We additionally demonstrate the strong performance of our model through extensive experiments and against numerous baselines.
5.1 Hue-shift MNIST Classification
We demonstrate improved generalization to global hue-shifts, and higher sample efficiency compared to a conventional CNN model with a similar number of parameters on our Hue-shift MNIST dataset. Our dataset is a variation of MNIST [32] with 60k training examples and 10k test examples classified into one of 10 categories. Additional details are provided in Appendix C.1.
(a) Z2CNN
(b) Hue-4
Generalization to global hue-shift.
The generalization performance of our hue-equivariant models, Hue-$N$ ($N$ indicates the cardinality of the discrete hue group), and a conventional CNN model, Z2CNN [6], are reported in Table 4. On the in-distribution test case ($A$/$A$), the performance of our model and the conventional CNN model are comparable. However, on the out-of-distribution test case ($A$/$B$), the performance of our model is preserved, while the performance of the conventional CNN model deteriorates significantly.
The performance gap on the out-of-distribution test case can be attributed to the difference in internal representation structure. To understand this better, we generate feature representation trajectories in the penultimate layer of our network and the baseline architecture by continuously varying the hue of an input image. We then visualize these trajectories using tSNE [33] projection (see Figure 5). For the baseline architecture, the trajectories of different digits overlap for some hues (Figure 5(a)), but for our model, the trajectories of different digits are confined to separate clusters (Figure 5(b)).
Model sample efficiency.
The sample efficiency of our hue-equivariant models (Hue-$N$) and a conventional CNN model (Z2CNN) are reported in Figure 4. The advantage of our hue-equivariant models increases as the percentage of the training set used decreases, illustrating improved sample efficiency [34].
(a) Hue-shifted feature maps
(b) Equivariance error
Figure 6: Hue equivariant feature maps. (a) We qualitatively compare hue-shifted feature maps produced using our lifting layer and the CEConv lifting layer. Highlighted feature maps obtained using our lifting layer (Hue-4) are qualitatively indistinguishable, while there are visible discrepancies in the feature maps of CEConv-4. (b) We quantitatively compare the equivariance-error of the hue-shifted feature maps for each approach. The equivariance error resulting from our lifting layer (Hue-4) is more than three orders of magnitude lower than the error resulting from CEConv-4.
5.2 Hue-shift 3D Shapes Classification
We demonstrate improved generalization to local hue-shifts, and significantly reduced equivariance error compared to CEConv on the 3D Shapes classification dataset [35]. The 3D Shapes dataset consists of RGB images of 3D shapes, where the color of the shape, the floor, and the walls vary across examples, as does the scale and orientation of the shape. There are 48k examples in the training set, and 12k examples in each of the test sets. Additional details are provided in Appendix C.2.
Table 2: Generalization to local hue-shift. Classification error on the 3D Shapes dataset is reported. Our model (Hue-$N$) and CEConv-$N$ achieve improved generalization performance over the conventional CNN model (Z2CNN). Our approach is robust to the choice of hue group discretization, while the performance of CEConv deteriorates when $N$ is not a symmetry of the axis-aligned RGB cube.
Network $A$/$A$ $A$/$B$ $A$/$C$ Params
Z2CNN 0.00 (0.00) 51.25 (9.59) 26.66 (19.60) 20,192
Hue-3* 0.00 (0.00) 0.03 (0.04) 0.04 (0.03) 19,478
@@ 5.2 Hue-shift 3D Shapes Classification
CEConv-3 0.00 (0.00) 0.02 (0.03) 0.05 (0.04) 26,441
CEConv-4 0.00 (0.03) 0.60 (0.25) 0.53 (0.43) 30,792
Generalization to local hue-shift.
The generalization performance of our hue-equivariant models, Hue-$N$ ($N$ indicates the cardinality of the discrete hue group), a conventional CNN model, Z2CNN [6], and the CEConv-$N$ models are reported in Table 2. Our Hue-4 model outperforms all other models on the in-distribution ($A/A$), global hue-shift out-of-distribution ($A/B$) and local hue-shift out-of-distribution ($A/C$) test sets. The performance of CEConv-4 is significantly worse than all other hue-equivariant models. We attribute the relatively poor generalization performance of the CEConv-4 model to the feature map equivariance error.
Lifting layer equivariance error.
The lifting layer equivariance error of our Hue-4 model and the CEConv-4 model are reported in Figure 6(b). As discussed in section 4, our lifting layer operates on HSL images and does not induce the invalid hue rotations observed in CEConv. We quantitatively assess how lifting layer invertibility impacts descriptor equivariance following [36]. We compute the equivariance error as the relative equivariance,
Equivariance $\text{Equivariance Error}=\frac{|f(\varphi_{h}(h_{i},x))-\phi_{h}(h_{i},f(x))|}{|f(\varphi_{h}(h_{i},x))+\phi_{h}(h_{i},f(x))|}$ (20)
and show that the equivariance error of CEConv-4 is more than three orders of magnitude higher than our Hue-4 model (see Figure 6(b)).
5.3 Camelyon17 Classification: Saturation shift in the wild
We demonstrate improved generalization to saturation-shifts compared to ResNet50 [3] and CEConv-3 on the Camelyon17 classification dataset [31]. The Camelyon17 dataset consist of images of human tissue collected from five different hospitals. Variation in tissue images results from variation in the data collection and processing procedures. The dataset consists of 387,490 examples, 302,436 of which are used for training and 85,054 of which are used for testing. Hospitals 1, 2, and 3 are represented in the training set, and hospital 5 is represented in the testing set. Additional details are provided in Appendix D.3.
Table 3: Generalization to saturation-shift. Classification error on the Camelyon17 dataset is reported. Our model (Sat-$N$ and Hue-$M$-Sat-$N$) achieves improved generalization performance over the conventional CNN model (ResNet50).
Networks Error Param
@@ 5.3 Camelyon17 Classification: Saturation shift in the wild
Sat-3* 16.08 (2.68) 23.3M
Hue-4-Sat-3* 19.06 (4.92) 23.0M
CEConv-3 28.76 (9.93) 23.1M
Generalization to saturation-shift.
We expand the notion of color equivariance proposed in [16] to capture variation in saturation. The generalization performance of our saturation-equivariant models, Resnet50 [3], and CEConv-3 are reported in Table 3. Our saturation- and color-equivariant models outperforms all other models, while CEConv performs comparably to the conventional CNN model.
5.4 Color shift in the wild
We demonstrate improved generalization to global color-shifts compared to ResNet [3] and CEConv on the Caltech-101 [37], Oxford-IIT Pets [38], Stanford Cars [39], STL-10 [40], CIFAR-10 and CIFAR-100 [41]. Additional details are provided in Appendix D.
Table 4: Generalization to color-shift. Classification error on the Caltech-101, Oxford-IIT Pets, Stanford Cars, CIFAR-10, CIFAR-100, STL-10 datasets are reported. Our models (Hue-$N$, Sat-$N$ and Hue-$M$-Sat-$N$) outperform CEConv-$N$ models on all datasets, and are competitive with ResNet.
Network Caltech 101 CIFAR-10 CIFAR-100 Stanford Cars Oxford Pets STL-10
ResNet 32.68 (1.55) 7.86 (1.14) 32.00 (0.63) 25.41 (0.96) 31.52 (2.05) 18.59 (1.65)
Hue-3* 34.32 (0.34) 8.41 (0.39) 33.28 (0.63) 22.19 (2.20) 31.37 (3.50) 19.82 (0.98)
Hue-4* 32.23 (1.07) 8.83 (0.64) 34.70 (0.89) 20.38 (1.06) 27.39 (0.68) 20.73 (1.11)
Hue-4-Sat-3* 38.14 (1.07) 10.68 (0.78) 33.27 (0.31) 24.79 (3.87) 29.84 (1.34) 20.53 (0.73)
Sat-3* 41.64 (1.40) 9.24 (0.27) 39.33 (0.45) 31.90 (10.03) 36.87 (5.57) 20.71 (1.10)
CEConv-3 34.74 (0.83) 8.86 (0.33) 34.95 (0.44) 23.97 (1.56) 31.08 (2.54) 24.29 (1.31)
CEConv-4 33.52 (0.48) 9.28 (0.24) 35.46 (0.35) 24.08 (0.66) 33.70 (1.50) 21.90 (1.64)
Generalization to color-shift.
The generalization performance of our hue-equivariant models, Hue-$N$, saturation-equivariant model, Sat-$N$, color-equivariant models Hue-$M$-Sat-$N$, ResNet44 model, and the CEConv-$N$ models are reported in Table 4.
Color sorting on CIFAR-10.
@@ 5.4 Color shift in the wild
While our hue-equivariant model performs on-par with the conventional CNN model on this dataset, the structure of our representation can be leveraged for tasks that are not possible with conventional architectures alone. To illustrate this, we use our model to sort images in the automobile class by hue (see Figure 7(a)). To produce an hue ordering, we compute the hue-shifted pairwise Euclidean distance between hue group feature maps of instances in the automobile class. Images $x_{1}$ and $x_{2}$ are close in the hue space if
$\arg\min_{i}\|\phi_{h}(h_{i},\Psi(x_{1})),\Psi(x_{2})\|=0,$ (21)
@@
where $\Psi(x_{j})$ denotes the penultimate feature map for image $j$ and $h_{i}$ is the $i$-th element of the hue group. Sorting using the penultimate feature map of the ResNet44 model is shown in Figure 7(b).
(a) Hue-$4$
(b) ResNet44
Figure 7: Color sorting on CIFAR-10. We show images from CIFAR-10 automobile class ordered by pairwise distance using Hue-4 and ResNet44 feature maps. The structure of the Hue-4 feature maps naturally allow for color based sorting, whereas the ResNet44 feature maps do not.
@@ 6 Conclusion
In this paper we address the challenge of learning hue- and saturation-equivariant representations. Leveraging the observation that these perceptual transformations have geometric structure, we propose a group structure for each, and a group convolutional neural network that is equivariant to the transformations by design. By encoding perceptual variation in a finite group structure, we achieve strong generalization performance and sample efficiency and can use learned representations for tasks such as color based sorting.
@@ 6.1 Limitations and future work
While our method is equivariant to hue and saturation by design, it is not guaranteed to be equivariant to luminance. However, since luminance is computed as a linear combination of the red, green and blue channel values of an input image, if the preprocessing statistics (e.g., data mean) on the training set are valid for the testing set, luminance is approximately unchanged. Additional limitations come from our choice of equivariant architecture. Group convolutional neural networks are more computationally expensive than their conventional counterparts since they require computation of the filter orbit at each layer. This is compounded with the requirement for a finite filter orbit which can at best approximate a continuous group. Future work will investigate a continuous representation of the group to address both of the aforementioned challenges.
7 Reproducibility Statement
Source code will be open-sourced upon acceptance. Proofs for hue and saturation group actions are shown in Appendix A. Proof of equivariance of the proposed lifting layer is shown in Section 4. Datasets, training and implementation details are described in Appendix C.
@@ 7 Reproducibility Statement
A Hue and saturation group action
@@ A.1 Hue group action
A.2 Saturation group action
@@
B Invalid hue rotation in CEConv
Figure 8: CEConv lifting. CEConv shift the hue of a filter in the RGB space by rotating its values about an axis passing through the point $p=(1,1,1)$. This approach results in invalid hue rotations for all discretizations of the hue group that are not symmetries of the axis-aligned RGB cube (i.e., $N=1$ and $N=3$). (Top) $N=3$. Rotations symmetric to axis-aligned RGB cube and yields valid rotations. (Bottom) $N=4$. Rotations yield results where invalid RGB values that lie outside of the RGB cube and needs to be projected.
C Datasets
C.1 Hue-shift MNIST
We introduce the Hue-shift MNIST dataset to evaluate the performance of our equivariant architectures in the presence of global hue shift. Examples in the training set are randomly assigned a hue between $0^{\circ}$ and $240^{\circ}$ (see Figure 9(a)) and examples in the test set are partitioned into an in-distribution set and an out-of-distribution set. In-distribution examples are randomly assigned a hue from the same distribution as the training set, and out-of-distribution examples are randomly assigned a hue outside of the training set distribution (i.e., between $240^{\circ}$ and $360^{\circ}$) (see Figure 9(b)).
(a) Dataset $A$ examples
(b) Dataset $B$ examples
(a) Dataset $A$ examples
(b) Dataset $B$ examples
(c) Dataset $C$ examples
C.2 Hue-shift 3D Shapes
We introduce the Hue-shift 3D Shapes dataset to evaluate the performance of our equivariant architectures in the presence of local hue shift. Examples in the training set are randomly assigned a hue from the first half of the color space (colors 0-4 as defined in Burgess and Kim [2018]) and examples in the test set are are partitioned into an in-distribution set and two out-of-distribution sets. The out-of-distribution test sets are designed to measure robustness to global hue-shift, and local hue-shift. In the out-of-distribution test set designed to measure robustness to global hue-shift, the color of the wall, the floor and the shape are randomly selected from the second half of the color space (colors 5-9 as defined in Burgess and Kim [2018]), and in the out-of-distribution test set designed to measure robustness to local hue-shift, the color of the wall and the floor are randomly selected from the first half of the color space, and the color of the shape is randomly selected from the second half of the color space.
D Implementation and training details
D.1 Hue-shift MNIST
We train our equivariant networks and the conventional architectures for 5 epochs with a batch size of 128. We optimize over a cross-entropy loss using the Adam optimizer Kingma and Ba [2014] with $\beta_{1}=0.9$, and $\beta_{2}=0.999$. We use an initial learning rate of $10^{-3}$ for the $\mathbb{Z}^{2}$ network, and $10^{-4}$ for the hue-equivariant network. We train CEConv architectures using the hyperparameters reported in Robert-Jan and van Gemert .
D.2 Hue-shift 3D Shapes
For this task we compare against the Z2CNN architectures proposed in Cohen and Welling [2016]. Our hue-equivariant architectures are designed with the same network structure as the baselines, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our hue-equivariant architectures, we perform hue group pooling to yield an hue-invariant representation.
Following Wiles et al. [2021], we train the ResNet architectures for 100k iterations with a batch size of 128. We optimize over a cross-entropy loss using SGD with a learning rate of $10^{-2}$. Images are down-sampled by a factor of 2 to train the CNN architectures. CNN architectures are trained as described in Section D.1.
D.3 Camelyon17
For this task, we compare against the ResNet50 architecture He et al. [2016]. Our hue- and saturation-equivariant architectures have the same number of layers as ResNet50, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our hue- and saturation-equivariant architectures, we perform a group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 10k iterations with a batch size of 32. We optimize over a cross-entropy loss using the Adam optimizer Kingma and Ba [2014] with an initial learning rate of $10^{-2}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. We train CEConv architectures using the hyperparameters reported in Robert-Jan and van Gemert .
Figure 11: Camelyon-17 dataset saturation statistics. Example images and saturation statistics by node (hospital).
We present experimental results for four discretizations of the saturation space. Our choice of discretizations, i.e., $d\in\{\nicefrac{{1}}{{20}},\nicefrac{{1}}{{10}},\nicefrac{{3}}{{20}},\nicefrac{{1}}{{2}}\}$, are determined with consideration of the saturation range of the training and validation set (see Figure 11). For a given choice of discretization $d$, an element $s_{i}\in S_{N}$ shifts the saturation of an input image by $s_{i}=i*d*c$, where $c$ is the maximum pixel value. We limit computational expense of lifting and convolution by performing these operations over the truncated set $\{s_{-1},s_{0},s_{1}\}$ for our Sat-3 model.
D.4 CIFAR-10
For this task, we compare against the ResNet44 architecture proposed in Cohen and Welling [2016]. Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet44, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 128. We optimize over a cross-entropy loss using SGD with an initial learning rate of $10^{-1}$ and a cosine-annealing scheduler. We evaluate the performance of our model in the presence of natural hue shifts on the CIFAR-100 Krizhevsky et al. [2009] classification dataset. The dataset consists of 60k examples, 50k of which are used for training and 10k for testing.
D.5 CIFAR-100
For this task, we compare against the ResNet44 architecture proposed in Cohen and Welling [2016]. Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet44, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 128. We optimize over a cross-entropy loss using SGD with an initial learning rate of $10^{-1}$ and a cosine-annealing scheduler. We evaluate the performance of our model in the presence of natural hue shifts on the CIFAR-100 Krizhevsky et al. [2009] classification dataset. The dataset consists of 60k examples, 40k of which are used for training, 10k for validation, and 10k for testing.
D.6 Caltech-101
For this task, we compare against the ResNet18 architecture proposed in He et al. [2016]. Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet18, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 16. We optimize over a cross-entropy loss using Adam Kingma and Ba [2014] with an initial learning rate of $10^{-2}$. We evaluate the performance of our model in the presence of natural hue shifts on the Caltech-101 Li et al. [2022] classification dataset. The dataset consists of 9,146 examples, $\nicefrac{{2}}{{3}}$ of which are used for training and $\nicefrac{{1}}{{3}}$ for testing.
D.7 STL-101
For this task, we compare against the ResNet18 architecture proposed in He et al. [2016]. Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet18, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 16. We optimize over a cross-entropy loss using Adam Kingma and Ba [2014] with an initial learning rate of $10^{-2}$. We evaluate the performance of our model in the presence of natural hue shifts on the STL-10 Coates et al. [2011] classification dataset. The dataset consists of 5k training examples and 8k testing examples.
D.8 Stanford Cars
For this task, we compare against the ResNet18 architecture proposed in He et al. [2016]. Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet18, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 16. We optimize over a cross-entropy loss using Adam Kingma and Ba [2014] with an initial learning rate of $10^{-2}$. We evaluate the performance of our model in the presence of natural hue shifts on the Stanford Cars Krause et al. [2013] classification dataset. As the original host for the Stanford cars dataset is no longer maintained, the dataset was constructed from the images Kaggle [2018], the devkit github [2024], and the annotation files Kaggle [2023]. The dataset consists of 8,144 training examples and 8,041 testing examples.
D.9 Oxford-IIT Pets
For this task, we compare against the ResNet18 architecture proposed in He et al. [2016]. Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet18, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 16. We optimize over a cross-entropy loss using Adam Kingma and Ba [2014] with an initial learning rate of $10^{-2}$. We evaluate the performance of our model in the presence of natural hue shifts on the Oxford-IIT Pets  Parkhi et al. classification dataset. The dataset consists of 3,680 training examples and 3,669 testing examples.